OpenStudy (trisaba):

what is the rule for the nth turn example a sub n =n for the set 1, 6, 11, 16, 21...

OpenStudy (trisaba):

the example was \[a _{n}=n\]

OpenStudy (trisaba):

@dan815 @zepdrix

zepdrix (zepdrix):

Hmm are you able to see a pattern in your numbers? :) Are we adding something each time? Multiplying by something each time? What type of change is happening from 1 to 6, from 6 to 11, and so on?

OpenStudy (trisaba):

adding 5

OpenStudy (trisaba):

that i knew the others i did not

zepdrix (zepdrix):

plus 5! Ok great. Let's come up with a rule for the second term, \(\large\rm a_2=a_1+5\) yes? How bout the third term, we would add 5 to the first term, twice, \(\large\rm a_3=1+5+5\) or we can write it this way, \(\large\rm a_3=1+5(2)\)

OpenStudy (trisaba):

but what about the first

zepdrix (zepdrix):

Well, we start at the first term, so we're adding none of those 5's, right?

zepdrix (zepdrix):

\[\large\rm a_1=1+5(0)\]

OpenStudy (trisaba):

so how about \[a _{n}=1+5(n-1)\]

zepdrix (zepdrix):

Mmm ya I like that idea! :) Maybe you noticed that the number of 5's we need is always `one less than our term`.\[\large\rm a_{\color{orangered}{2}}=1+5(\color{orangered}{1})\]\[\large\rm a_{\color{orangered}{3}}=1+5(\color{orangered}{2})\]\[\large\rm a_{\color{orangered}{4}}=1+5(\color{orangered}{3})\]\[\large\rm a_{\color{orangered}{n}}=1+5(\color{orangered}{n-1})\]So yes :) gj

OpenStudy (trisaba):

what if the set was 1, 2, 4, 8, 16 you would multiply by 2

OpenStudy (trisaba):

@phi

OpenStudy (trisaba):

@zepdrix

OpenStudy (phi):

yes, you multiply the "previous value" to get the next value

OpenStudy (trisaba):

what is the rule

OpenStudy (phi):

or , for n (the term) =1,2,3,4, the corresponding value is 2^(n-1)

OpenStudy (trisaba):

thanks

OpenStudy (phi):

notice (ignoring the 1 for the moment) your numbers 1,2,4,8,16 can be written 1, 2, 2*2, 2*2*2, 2*2*2*2

OpenStudy (phi):

if you know 2^0 is 1 and 2 is the same as 2^1 and 2*2 is 2^2 you see the pattern 2^0, 2^1, 2^2, 2^3, 2^4 and if we count the terms using n=1,2,3,4, the nth term is 2^(n-1)

OpenStudy (phi):

For example, the first term (when n=1) gives the answer 2^(1-1) = 2^0 = 1

OpenStudy (trisaba):

true what if the signs swap like -4, 8, -12, 16, -20 what happens to the rule because you have to multiply and add sorry for asking to much

OpenStudy (phi):

one way is to first ignore the signs, so you have 4,8,12,16,20 can you write a formula for then nth term ?

OpenStudy (trisaba):

\[a _{n}=4n\]

OpenStudy (phi):

if you use the formula 4n with n=1,2,3,etc we get 4, 8, 12, 16, etc that looks good to get alternating minus signs we use a trick (-1)^n when n is odd gives us -1 (-1)^n when n is even gives us +1 so multiply your formula 4n by (-1)^n (-1)^n 4n

OpenStudy (trisaba):

\[a _{n}=(-1)^{4n}\]

OpenStudy (phi):

\[ a_n= (-1)^n 4 n \]

OpenStudy (trisaba):

got it

OpenStudy (trisaba):

some are though unfamiliar those were easy to pop out

OpenStudy (trisaba):

this one is unfamiliar how do you do that there is niether a common difference or common ratio 2, 9, 28, 65, n=5

OpenStudy (phi):

does the n=5 mean they want the 5th term?

OpenStudy (trisaba):

need rule

OpenStudy (trisaba):

and 5th

OpenStudy (phi):

you can find differences of differences like this 2, 9, 28, 65 7 19 37 12 18 6 we have to assume that the 6 represents the pattern thus, to extend the pattern we do 2, 9, 28, 65 7 19 37 12 18 6 6 the next row above the 6 will be 18+6 2, 9, 28, 65 7 19 37 12 18 24 6 6 extending the row above the 24, we get 37+24 2, 9, 28, 65 7 19 37 61 12 18 24 6 6 finally, the top row is extended by 65+61 2, 9, 28, 65 126 7 19 37 61 12 18 24 6 6

OpenStudy (trisaba):

desrcibe and write the rule

OpenStudy (dan815):

its ++5 every time so

OpenStudy (dan815):

after 5 times it will be plus n*5

OpenStudy (trisaba):

no the new one not the 5

OpenStudy (trisaba):

2 9 28 65 i need n=5 and rule

OpenStudy (dan815):

http://prntscr.com/a7kgai

OpenStudy (trisaba):

i dont get the rule though

OpenStudy (trisaba):

i dont get the rule though